# Less is more: Probabilistic reduction is best explained by small-scale predictability measures

### v1

@@ Abstract
The primary research questions of this paper center on defining the amount of context that is necessary and/or appropriate when investigating the relationship between language model probabilities and cognitive phenomena. We investigate whether whole utterances are necessary to observe probabilistic reduction and demonstrate that n-gram representations suffice as cognitive units of planning.
Less is more: Probabilistic reduction is best explained by small-scale predictability measures
Cassandra L. Jacobs Department of Linguistics University at Buffalo Buffalo, NY, USA cxjacobs@buffalo.edu          Andrés Buxó-Lugo Department of Psychology University at Buffalo Buffalo, NY, USA buxolugo@buffalo.edu
Anna K. Taylor EURECOM Université Paris - Sorbonne anna.taylor@eurecom.fr          Marie Leopold-Hooke Department of Computer Science and Engineering University at Buffalo Buffalo, NY, USA marieleo@buffalo.edu
@@ 1 Introduction
The articulatory dynamics of language production are sensitive to linguistic statistics, such as the frequencies of words, the phrases that those words appear in, and the relative probability of those words given the context dammalapati2021effects; bell2009predictability; harmon2021theory. Familiar words and phrases are assumed to be shorter due to factors like motor practice bybee2006usage, or other factors like the general ease of retrieval arnold2012audience. Previous studies have demonstrated that predictability influences articulatory dynamics, such that the more probable a word is, the shorter in duration the words tend to be lieberman1963some.
@@
Words can be predictable in different ways. We can both anticipate the future, (e.g., "carve a…" is likely to end with pumpkin) as well as confirm that a word is supported by later information (e.g., pumpkin pie). Whereas the influence of word frequency on reduction could be explained by motor practice, findings that words are reduced when they are predictable has been argued to support rational production models. Conversely, lengthening improbable words is believed to afford listeners extra time to interpret the speech signal and minimize the odds of miscommunication aylett2004smooth; degen2020redundancy; futrell2023information; mahowald2013info; clark_relationship_2025; piantadosi2011word. However, probabilistic reduction effects are apparently illusory or contradictory depending on the level of granularity at which probabilities are computed (n-gram versus neural language model), and the type of word (content vs. function; bell2009predictability). Recent work has argued that large language models are uniquely suited to simultaneously explain probabilistic reduction and repetition reduction upadhye2025back; clark_relationship_2025. These studies have argued that n-grams are fundamentally unable to account for enough context to reliably model reduction relative to LLMs, which can encode up to 1024 tokens of context.
@@
However, the cognitive plausibility of large language models as applied toward language production is unclear. Firstly, it is unknown what knowledge speakers draw on during language production, and how a speaker might efficiently estimate a word’s probability in context. The broad consensus in language production research is that speakers do not plan their utterances entirely in advance, and instead plan their utterances relatively incrementally ferreira2002incremental; futrell2023information; guhe2020incremental; pickering2013forward. So, the language production system might plan utterances at a lower level of granularity (e.g., one or two nearby words), or perhaps a larger one momma2021filling. Here we show that local probabilities suffice for the computation of probabilistic reduction, particularly when considering the computational efficiency of phrase-level retrieval arnon2013more; bannard2008stored.
@@
To assess this possibility, the experiments described in this paper compare and contrast probability estimates of words derived from n-gram and LLM estimates, which are then incorporated as predictors in regression analyses of whole-word phonetic duration. We analyze four spontaneous speech corpora of American English: the Buckeye Corpus pitt2005buckeye, the Corpus of Regional African American Language (CORAAL; kendall2013speech; kendall2018corpus), Switchboard-NXT calhoun2010nxt of upadhye2025back, and the CANDOR corpus reece2023candor from clark_relationship_2025. Critically, our analyses include critical prosodic covariates that are known to affect word durations bishop2018anticipatory; baker2009variability (Section §2) and which we show to be partially confounded with language model probabilities (Section §5).
@@
We demonstrate the utility of small-context modeling in Section §4 by showing that LLM probabilities extracted over intonational units, rather than whole conversational turns, better account for the effects of predictability on duration. We then show in Section §5 and §6 that statistics from large language models are not necessary to observe probabilistic reduction. In fact, the results of these analyses are broadly consistent with each other that LLM predictors produce worse model fits to durations than n-gram probabilities. We then examine the regression models across all datasets and observe that now-common practices for model comparison have masked the extremely small effect size of probabilistic factors on phonetic duration (as discussed in bell2009predictability).
@@ 2 Prosodic Preliminaries
A word’s duration (and, indeed, other dimensions of speech prosody) are sensitive to structural, probabilistic, and utterance-specific factors baker2009variability. In order to accurately gauge the contribution of probabilistic factors to duration, and determine whether large language models or multiword representations are more appropriate, we discuss these critical covariates.
@@ 2.1 Structural and lexical influences on duration
“Structural” factors that are known to affect duration include pauses, syntactic or prosodic phrase boundaries byrd2003elastic; klatt1975vowel; turk2007multiple; wightman1992segmental, in addition to lexical information such as number of phonemes and/or syllables, a word’s phonotactic probability goldrick2008phonotactic, and its phonological neighborhood density gahl2012reduce; gahl2016many.
@@ 2.2 Utterance length and position confounds on duration
Word durations have been shown to vary based on utterance planning/production dynamics. For example, words produced in isolation tend to have the longest durations as speakers may be disfluent and using timing information to hold the floor in a conversation clark2002using. Occasionally, researchers incorporate pause duration or pause location as a control variable clark_relationship_2025. This work instead conceptualizes phrase-final lengthening as a relative position effect, such that words closer to the end of a sentence tend to be longer.
Additionally, “anticipatory shortening” is a pattern where the durations of a syllable in a phrase are negatively correlated with the length of the phrase bishop2018anticipatory. This has been argued to reflect scope of planning effects, as syllable durations throughout the utterance are affected by a factor that is not made explicit until the production is complete. Thus, particularly for longer utterances, speech rates may be planned out somewhat in advance based on the approximate number of words or syllables a speaker plans. In contrast to prior work which has used speaker speech rate as a covariate, we use utterance-specific phrase length to account for gross timing patterns.
We determine utterances boundaries based on inter-pause units, which produce prosodic phrases, which have been argued to represent a fundamental unit of planning in production wheeldon2002minimal. Segmenting the utterance along inter-pause intervals and encoding a word’s relative position in an utterance produces the canonical anticipatory shortening effects across all datasets except Switchboard, which we show in Figure 1. All lengths of utterance show the standard phrase-final lengthening effect.
@@ 2.3 Probabilistic factors
“Probabilistic” factors that affect reduction include a word’s frequency, its forward predictability, and its backward predictability clark_relationship_2025; bell2009predictability; upadhye2025back. Words that appear in frequent phrases are also typically reduced arnon2013more. In all cases, a word is expected to be reduced in duration when it is more probable in context. Additionally, a word’s informativity has a moderating effect, such that words that are typically unpredictable will tend to be less reduced seyfarth2014word.
Taking bigram-based conditional probabilities as an example, a word’s probability at position $i$ in context can be broadly generalized as $\frac{p(w_{i-1},w_{i})}{p(w_{i})}$ for forward transition probability, and $\frac{p(w_{i},w_{i+1})}{p(w_{i+1})}$ for backward transition probabilities. Large language models typically incorporate more context than n-gram probabilities, and can be computed for very large sequences of text, such as an entire conversation clark_relationship_2025, but the computation is similar.
@@
Table 1: Expected relationships of predictors on whole-word duration.
@@ 2.4 Covariate selection
To control for non-probabilistic factors that could influence phonetic duration, we extracted several predictor variables from previously established norming databases. For consistency with previous work, we use frequency information from the SUBTLEX-US corpus brysbaert2009moving. We obtain norms for a word’s length in number of phonemes from chee2020consistency. Number of syllables was not found to be a significant predictor in any analysis, and was recommended for exclusion in all model comparisons; therefore, we do not report any models with it included. Each word’s forward and backward transition probability is computed per dataset by the relevant n-gram or neural language model, described in each section. Table 1 shows the predicted outcomes from these analyses.
To account for phrase-final lengthening, we compute a transform on relative position, itself defined as the position ii111ii is strictly positive and non-zero. for a word $w_{i}$ out of all of the $n$ words in the sentence. We compute an inverse position value that produces higher values the closer the sentence is to the end.
@@
At the end of the sentence, this value is undefined; therefore, we impute the maximum value from other non-initial positions in the other sentences in each corpus to determine this inverse position value.
@@
We include random intercepts for each word in a particular part of speech (POS) category (e.g., extract as a noun or verb) to account for variation that might arise for homographic words or words used in different sentences as well as random intercepts by talker. For analysis of Buckeye and CORAAL in Section §6, course-grained POS tags were obtained through spaCy Honnibal_spaCy_Industrial-strength_Natural_2020. We relied on the Penn Treebank POS tags in Section §5, and no POS labels were available for Section §4.
3 Cognitive plausibility of backward language modeling
For speakers to know that a particular word will be easily understood because it is partly redundant with the downstream context, and can therefore be reduced aylett2004smooth; jaeger2010redundancy, the language production system must be able to compute such a probability efficiently.
The cognitive plausibility of LLMs relative to n-grams is less than obvious. It is widely believed that speakers store multiword sequences in long-term memory and that doing so promotes production fluency arnon2013more; bannard2008stored; morgan2024productive; siyanova2018production. By contrast, computing probabilities from a language model typically involves encoding the full, known upstream and/or downstream context to estimate a word’s probability. Alternately, some have applied the fill-in-the-middle procedure of bavarian2022efficienttraininglanguagemodels to more efficiently extract words’ probabilities, but this procedure does not clearly correspond to any known cognitive processes that speakers engage in to learn language, nor to perform fluent production. So, while the estimates from such a model may produce higher-quality predictors, the cognitive plausibility of these predictors is suspect. Finally, as upadhye2025back point out, solutions built on beam search are too expensive to be plausible accounts of production either meister-etal-2020-beam; pmlr-v97-cohen19a.
We propose that one efficient way for predictability to influence durations could be through the retrieval of multiword sequences (e.g., trigrams) from memory bannard2008stored; siyanova2018production. The subsequent experiments establish the feasibility of retrieving multiword sequences from memory as a way to explain probabilistic reduction.
@@ 3 Cognitive plausibility of backward language modeling
4 Establishing short-range probabilistic reduction in clark_relationship_2025
@@
In their analysis of the CANDOR corpus, clark_relationship_2025 found that words that are less predictable from the upstream context are produced with longer durations. As clark_relationship_2025 point out, the general length of a turn in conversation is rarely very long, typically under 50 words. Thus, their use of GPT-2’s full context window (up to 1024 tokens) to encode the entire turn is likely overly powerful for estimating a word’s probability. Thus, in this section, we re-analyzed the CANDOR corpus, but extracted language model probabilities only over intonational phrases using the probability estimation method of pimentel_how_2024.
As part of this reanalysis, we re-tokenized the utterances along pause boundaries to produce intonational phrases, which shortens the typical lengths of utterances substantially. We expect that smaller amounts of context will produce better estimates of a word’s probability in context, and therefore better predict the words’ durations.
@@
We find that intonational phrase-level probabilities are more robust predictors of probabilistic production. Moreover, the two probability measures are highly correlated (Spearman $\rho$ = 0.69), and including both the short-term and long-term predictability measures hinders model convergence. Critically, the effect of predictability on duration is fully subsumed by the small-context estimate. The results of the regression model are included in Table 2. We visualize the stronger relationship to phonetic duration in Figure 2.
@@
(Intercept) 0.254 0.034 7.455 <.001
SUBTLEX word frequency log -0.167 0.005 -28.980 <.001
Number of phonemes 0.142 0.003 55.654 <.001
Utterance length in words ($n$) log -0.055 0.003 -20.557 <.001
Short-range probability log -0.021 0.001 -38.126 <.001
Relative position inverse log 0.188 0.001 128.406 <.001
@@
Figure 2: Correlation between log language model probability and words’ durations (log ms) for short- and long-range language model probabilities.
@@
5 N-gram versus language model predictability and probabilistic reduction in upadhye2025back

### v2

@@ Abstract
We explore the cognitive representations and mechanisms that underpin probabilistic reduction in phonetic duration. Recently, massive statistical models of language have been developed that capture long-distance dependencies between words, but the plausibility of these systems as models of human cognition is unclear. Following up on two recent studies, we test in three studies whether the amount of context that these systems use is necessary and/or appropriate when investigating the relationship between language model probabilities and phonetic duration. We establish that the retrieval of multi-word sequences suffice as cognitive units of planning and that retrieval of frequent sequences promotes reduction.
keywords:
language production , probabilistic reduction , phonetic duration , prosody , large language models , cognitive plausibility , phrase representations
journal: In submission to the Journal of Memory and Language
[…]
[UBL]organization=Department of Linguistics, University at Buffalo, city=Buffalo, state=NY, country=USA
[…]
[UBP]organization=Department of Psychology, University at Buffalo, city=Buffalo, state=NY, country=USA
[…]
[EURECOM]organization=EURECOM, city=Biot, country=France
[…]
[PARIS]organization=Université Paris - Sorbonne, city=Paris, country=France
[…]
[UBC]organization=Department of Computer Science and Engineering, University at Buffalo, city=Buffalo, state=NY, country=USA
@@ 1 Introduction
The articulatory dynamics of language production are sensitive to linguistic statistics, such as the frequencies of words, the phrases that those words appear in, and the relative probability of those words given the context (dammalapati2021effects; bell2009predictability; harmon2021theory). Familiar words and phrases are assumed to be shorter due to factors like motor practice (bybee2006usage), or other factors like the general ease of retrieval (arnold2012audience). Previous studies have demonstrated that predictability influences articulatory dynamics, such that the more probable a word is, the shorter in duration the words tend to be (lieberman1963some).
@@
Words can be predictable in different ways. We can both anticipate the future, (e.g., “carve a…” is likely to end with pumpkin) as well as confirm that a word is supported by later information (e.g., pumpkin pie). Whereas the influence of word frequency on reduction could be explained by motor practice, findings that words are reduced when they are predictable has been argued to support rational production models. Conversely, lengthening improbable words is believed to afford listeners extra time to interpret the speech signal and minimize the odds of miscommunication (aylett2004smooth; degen2020redundancy; futrell2023information; mahowald2013info; clark_relationship_2025; piantadosi2011word). However, probabilistic reduction effects are apparently illusory or contradictory depending on the level of granularity at which probabilities are computed (n-gram versus neural language model), and the type of word (content vs. function; bell2009predictability). Recent work has argued that large language models are uniquely suited to simultaneously explain probabilistic reduction and repetition reduction (upadhye2025back; clark_relationship_2025).
These studies have argued that n-grams are fundamentally unable to account for enough context to reliably model reduction relative to large language models (LLMs), which can encode very long sequences of prior context.
However, the cognitive plausibility of large language models as applied toward language production is unclear. Firstly, it is unknown what knowledge speakers draw on during language production, and how a speaker might efficiently estimate a word’s probability in context. For speakers to know that a particular word will be easily understood because it is partly redundant with the downstream context, and can therefore be reduced (aylett2004smooth; jaeger2010redundancy), the language production system must be able to compute such a probability efficiently. Computing probabilities from a language model typically involves encoding the full, known upstream and/or downstream context to estimate a word’s probability. Alternately, some have applied the fill-in-the-middle procedure of bavarian2022efficienttraininglanguagemodels to more efficiently extract words’ probabilities, but this procedure does not clearly correspond to any known cognitive processes that speakers engage in to learn language, nor to perform fluent production. Additionally, as upadhye2025back point out, solutions built on beam search or sampling (giulianelli2024generalized) are too expensive to be plausible accounts of production either (meister-etal-2020-beam; pmlr-v97-cohen19a).
@@
Long-distance context sensitivity of the process of multiword production is at odds with the broad consensus in language production research that speakers do not plan their utterances entirely in advance, and instead plan their utterances relatively incrementally (ferreira2002incremental; futrell2023information; guhe2020incremental; meyer1996lexical; pickering2013forward). So, the language production system might plan utterances at a lower level of granularity (e.g., one or two nearby words), or perhaps a larger one when certain kinds of long-distance dependencies are considered (momma2021filling). It is widely believed that speakers store multiword sequences in long-term memory and that doing so promotes production fluency (arnon2013more; bannard2008stored; morgan2024productive; siyanova2018production). Here we show that local probabilities suffice for the computation of probabilistic reduction, and phrase-based representations are more plausible units of retrieval when considering the computational efficiency of phrase-level retrieval.
@@
To assess these competing accounts, the experiments described in this paper compare and contrast probability estimates of words derived from n-gram and LLM estimates, which are then incorporated as predictors in regression analyses of whole-word phonetic duration. We analyze four spontaneous speech corpora of American English: the Buckeye Corpus (pitt2005buckeye), the Corpus of Regional African American Language (CORAAL; kendall2013speech; kendall2018corpus), Switchboard-NXT (calhoun2010nxt) of upadhye2025back, and the CANDOR corpus (reece2023candor) from clark_relationship_2025. Critically, our analyses include important utterance-level prosodic covariates that are known to affect word durations (bishop2018anticipatory; baker2009variability; Section §2).
@@
We demonstrate the utility of small-context modeling in Section §3 by showing that LLM probabilities extracted over intonational units, rather than whole conversational turns, better account for the effects of predictability on duration. We then discuss that language model probabilities may be partially confounded with language model probabilities (Section §4). We additionally show in Section §4 and §5 that statistics from large language models are not necessary to observe probabilistic reduction. In fact, the results of n-gram analyses are broadly consistent and demonstrate that LLM predictors produce worse model fits to durations than n-gram probabilities. In Section 6 we then examine the regression models across all datasets and observe that now-common practices for model comparison have masked the extremely small and potentially unstable effect size of probabilistic factors on phonetic duration (as discussed in bell2009predictability). We ultimately provide a verbal account building on multi-word linguistic representations that can explain probabilistic reduction efficiently.
@@ 2 Prosodic Preliminaries
Words’ durations (of course in addition to other dimensions of speech prosody) are sensitive to structural, probabilistic, and utterance-specific factors (baker2009variability). In order to accurately gauge the contribution of probabilistic factors to duration, and determine whether large language models or multiword representations are more appropriate, we discuss these critical covariates.
@@ 2.1 Structural and lexical influences on duration
“Structural” factors that are known to affect duration include pauses, syntactic or prosodic phrase boundaries (byrd2003elastic; klatt1975vowel; turk2007multiple; wightman1992segmental), in addition to lexical information such as number of phonemes and/or syllables (TurkAliceE.1997Tdoa), a word’s phonotactic probability (goldrick2008phonotactic), and its phonological neighborhood density (gahl2012reduce; gahl2016many).
@@ 2.2 Utterance length and position confounds on duration
Word durations have been shown to vary based on utterance planning/production dynamics. For example, words produced in isolation in spontaneous speech tend to have the longest durations as speakers may be disfluent and use timing information to hold the floor in a conversation (clark2002using). Occasionally, researchers incorporate pause duration or pause location as a control variable (clark_relationship_2025). This work instead conceptualizes phrase-final lengthening as a relative position effect, such that words closer to the end of a sentence tend to be longer.
Relatedly, “anticipatory shortening” is a pattern where the durations of a syllable in a phrase are negatively correlated with the length of the phrase (bishop2018anticipatory). This has been argued to reflect scope of planning effects, as syllable durations throughout the utterance are affected by a factor that is not made explicit until the production is complete. Thus, particularly for longer utterances, speech rates may be planned out somewhat in advance based on the approximate number of words or syllables a speaker plans. Similar findings have been demonstrated in fundamental frequency, such that longer utterances typically start with higher pitch than shorter utterances, suggesting that producers are broadly aware of how long they want to speak (liberman1977stress). In contrast to prior work which has used speaker speech rate (upadhye2025back) or numeric word position as a covariate (clark_relationship_2025), we use utterance-specific phrase length to account for gross timing patterns.
To better model utterance length, we chunk utterance boundaries into inter-pause units based on human-annotated pause boundaries. The resulting prosodic or intonational phrases have been argued to represent a fundamental unit of planning in production (wheeldon2002minimal). Segmenting speech along inter-pause intervals and encoding a word’s relative position in an utterance produces the canonical anticipatory shortening effects across all datasets except Switchboard, which we show in Figure 1. All lengths of utterance show the standard phrase-final lengthening effect.
@@ 2.3 Probabilistic factors
“Probabilistic” factors that affect reduction include a word’s frequency, its forward predictability, and its backward predictability (clark_relationship_2025; bell2009predictability; upadhye2025back). Words that appear in frequent phrases are also typically reduced (arnon2013more). In all cases, a word is expected to be reduced in duration when it is more probable in context. Additionally, a word’s informativity has a moderating effect, such that words that are typically unpredictable will tend to be less reduced (seyfarth2014word).
In the case of two-word sequences, or bigrams, the probability of a word $w$ at position $i$ in context $c$ can be broadly generalized as
$c=w_{i-1},\frac{p(w_{i-1},w_{i})}{p(w_{i})}$ (1)
for the forward transition probability of encountering the word $w_{i}$ next, and
$c=w_{i+1},\frac{p(w_{i},w_{i+1})}{p(w_{i+1})}$ (2)
for the backward transition probability of having previously encountered $w_{i}$ given the next word.
Large language models typically incorporate more context than n-gram probabilities, and can be computed for very large sequences of text, such as an entire conversation (clark_relationship_2025), but the idea behind the computation is similar. For language models (including n-grams) with arbitrary memory, instead of conditioning on the immediately preceding word, a word’s probability can depend on the complete chain of preceding words in the following way:
$p(w_{i}|c)\propto p(w_{i}|w_{1}...w_{i-1})$ (3)
The probability estimated in Equation 3 is intractable to compute in practice, particularly for longer sequences. The application of autoregressive language models (e.g., GPT-2; radford_language_2019) within psycholinguistics has provided researchers with a means to quantify a word’s probability of being the next word given the preceding context through a transformation of the language model’s representation of the prior context. In practice, these probabilities are computed by multiplying the last hidden state with a matrix that produces a vector of activations, which is then scaled to produce a quasi-probability distribution with a softmax (historically sigmoid) transformation (rumelhart1986general).
@@
Table 1: Expected relationships of critical predictors on whole-word duration.
@@ 2.4 Covariate selection
To control for non-probabilistic factors that could influence phonetic duration, we extracted several predictor variables from previously established norming databases. For consistency with previous work, we use frequency information from the SUBTLEX-US corpus (brysbaert2009moving). We obtain norms for a word’s length in number of phonemes from chee2020consistency. Number of syllables was not found to be a significant predictor in any analysis, and was recommended for exclusion in all model comparisons; therefore, we do not report any models with it included. Such a finding may be due to the existence of other covariates in the model, particularly by-word random intercepts and other predictors in the fixed effects that encode length, as syllable structure is known to influence word duration (TurkAliceE.1997Tdoa). Each word’s forward and backward transition probability is computed per dataset by the relevant n-gram or neural language model, described in each section. Table 1 shows the predicted outcomes from these analyses.
To account for phrase-final lengthening, we compute a transform on relative position, itself defined as the position $i$ for a word $w_{i}$ out of all of the $n$ words in the sentence. $i$ is strictly positive and non-zero. We compute an inverse position value that produces higher values the closer the sentence is to the end.
@@
At the end of the sentence ($i=n$), this value is undefined; therefore, we impute the maximum value from other non-initial positions in the other sentences in each corpus to determine this inverse position value.
@@
Where possible, we include random intercepts for each word in a particular part of speech (POS) category (e.g., extract as a noun or verb) to account for variation that might arise for homographic words or words used in different sentences as well as random intercepts by talker. For analysis of Buckeye and CORAAL in Section §5, course-grained POS tags were obtained through spaCy (Honnibal_spaCy_Industrial-strength_Natural_2020). We relied on the Penn Treebank POS tags in Section §4, and no POS labels were available for Section §3.
2.5 Data availability
All data and analyses are available on the Open Science Foundation (external link).
@@ 2.5 Data availability
3 Study 1: Establishing short-range probabilistic reduction in clark_relationship_2025
If speakers modulate word durations to make unpredictable words longer and predictable words shorter, this would suggest that speakers are taking into account the needs of the listener. Some long-distance dependencies, such as a word’s discourse status (e.g., given versus new; fowler1987talkers), influence word durations as well. In their analysis of the CANDOR corpus, clark_relationship_2025 computed a transformation of next-word probabilities known as surprisal and found that words that were less predictable from the upstream context were produced with longer durations. They reasoned that the long-distance abilities of the large language model allowed them to account not just for the predictability effects of words at short distances, but also at long distances.
@@ 3 Study 1: Establishing short-range probabilistic reduction in clark_relationship_2025
As clark_relationship_2025 point out, the general length of a turn in conversation is rarely very long, typically under 50 words. Thus, their use of GPT-2’s full context window (up to 1024 tokens) to encode the entire turn is likely overly powerful for estimating a word’s probability. We expect that smaller amounts of context will produce better estimates of a word’s probability in context, and therefore better account for variability in word durations. Thus, in this section, we re-analyzed the CANDOR corpus, but extracted language model probabilities using the probability estimation method of pimentel_how_2024 only over intonational phrases (wheeldon2002minimal). To do so, we re-tokenized CANDOR utterances along known pause boundaries to produce intonational phrases, which shortens the typical lengths of utterances substantially. As the model estimates were extracted over the same sentences, the two probability measures are highly correlated (Spearman $\rho$ = 0.69).
@@
We find that intonational phrase-level probabilities are more robust predictors of probabilistic reduction. Critically, the effect of predictability on duration is fully subsumed by the small-context estimate – there is no contribution to model fit of long-distance probabilities. The results of the regression model are included in Table 2. We visualize the stronger relationship to phonetic duration in Figure 2.
@@
(Intercept) 0.254 0.034 7.455 ¡.001
SUBTLEX word frequency log -0.167 0.005 -28.980 ¡.001
Number of phonemes 0.142 0.003 55.654 ¡.001
Utterance length in words ($n$) log -0.055 0.003 -20.557 ¡.001
Short-range probability log -0.021 0.001 -38.126 ¡.001
Relative position inverse log 0.188 0.001 128.406 ¡.001
@@
Figure 2: Correlation between log language model probability and words’ durations (log ms) for short- and long-range language model probabilities. X axis is binned for visualization.
@@
4 Study 2: N-gram versus language model predictability and probabilistic reduction in upadhye2025back